OpenStudy (anonymous):

Solve 4(c - 3) = 4c + 8. Answer c = 11 c = 5 c = 0 No Solution

OpenStudy (anonymous):

4(c - 3) = 4c + 8 4c - 12 =4c + 8 4c = 4c + 20 in the end the c's cancel out, so the answer is No solution.